# Sample Compression Scheme Reductions

### v1

@@ 3.1 Additional Assumption: Existence of Proper or Majority Vote Binary Compression
Suppose any binary concept class $\mathcal{C}$ with VC dimension $\mathrm{VC}<\infty$ has a compression scheme of size $f(\mathrm{VC})$ which is a proper compression scheme or a majority vote compression scheme (see Definition 3.4). Then any multiclass concept class (label set is allowed to be of infinite size) with a finite graph dimension $\mathrm{d}_{\mathrm{G}}<\infty$ admits a compression scheme of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}})\right)$.
In particular, we recover the best known bounded-size multiclass compression scheme of size $2^{O\mathopen{}\left(\mathrm{d}_{\mathrm{G}}\right)}$ [David et al., 2016], via a reduction to binary compression scheme.
@@
We now construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given realizable $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, let $T=\{((x_{i},y_{i}),1):i\in[n]\}$. We let our compression $\kappa(S)$ return the $(x,y)$ pairs from $\kappa_{b}(T)$, which have size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}})\right)$. Reconstruction: We have that $\kappa_{b}(T)=\{((x,y),1):(x,y)\in\kappa(S)\}$, so we can immediately recover $\kappa_{b}(T)$. We know that $\rho_{b}(\kappa_{b}(T))$ is correct on all possible $T$. Additionally, since we predict $1$ for at most one $y$ for each $x$, we can conclude that $\rho_{b}(\kappa_{b}(T))$ is correct on all of $S_{\mathcal{Y}}$, as desired. Note that if the reconstruction $\rho_{b}$ is proper, then it will output exactly one $1$ for each $x$. Additionally, it the reconstruction is a majority of learners from the class, where we break ties to favor predicting 00, then it will output at most one $1$ for each $x$.
@@ 3.2 Additional Assumption: Existence of Stable Binary Compression
Suppose that for binary concept classes with finite VC dimension $\mathrm{VC}<\infty$, there exists a stable sample compression scheme of size $f(\mathrm{VC})$. Then, for multiclass concept classes with a finite graph dimension $\mathrm{d}_{\mathrm{G}}<\infty$ and finite label space, there exists a stable sample compression scheme of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}})\right)$.
@@
i.e., it is an inflated $\kappa_{b}(S)$ to include all the labels for each $x$. Since the compression scheme is stable, and since $\kappa_{b}(S_{\mathcal{Y}})\subseteq\kappa_{b}^{\prime}(S_{\mathcal{Y}})\subseteq S_{\mathcal{Y}}$, we must have that $\kappa_{b}(\kappa_{b}^{\prime}(S_{\mathcal{Y}}))=\kappa_{b}(S_{\mathcal{Y}})$. We can let $\kappa(S)$ return the $(x_{i},y_{i})$ pairs such that $x_{i}$ is a member of $\kappa_{b}(S_{\mathcal{Y}})$. This will have size $O(f(\mathrm{d}_{\mathrm{G}}))$. Reconstruction: Given $\kappa(S)$, we can reconstruct $\kappa^{\prime}_{b}(S_{\mathcal{Y}})$ by inflating the dataset to include all labels for each $x$. Then, we can apply $\kappa_{b}$ to get $\kappa_{b}(S_{\mathcal{Y}})$ (using the fact that it is a stable compression scheme), and then apply $\rho_{b}$ to get the desired compression scheme.
@@
Consider the class $\mathcal{C}\subseteq\mathbb{Y}^{\mathcal{X}}$, where $\mathcal{X}=\mathcal{Y}=\mathbb{R}$, of two piecewise thresholds, i.e. $\{g_{t,y_{1},y_{2}}:t,y_{1},y_{2}\in\mathbb{R}\}$ where
@@ 3.3 Agnostic Multiclass Sample Compression Scheme
that for binary classes with VC dimension $\mathrm{VC}<\infty$, there exists a sample compression scheme of size $f(\mathrm{VC})$. Then, for a multiclass class with a finite graph dimension $\mathrm{d}_{\mathrm{G}}<\infty$ and a finite label set $|\mathcal{Y}|$, there exists an agnostic sample compression scheme of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}})\log|\mathcal{Y}|\right)$.
@@ 4 Compression for Regression
Suppose that for binary classes with VC dimension $\mathrm{VC}<\infty$, there exists a sample compression scheme of size $f(\mathrm{VC})$. Then, for a $[0,1]$-real-valued class with pseudo-dimension $d_{\mathrm{P}}<\infty$, there is an $\epsilon$-approximate compression scheme with respect to the $\ell_{\infty}$ loss with size $O(f(d_{\mathrm{P}})\log(\frac{1}{\epsilon}))$. Furthermore, for $p\in[1,\infty)$, there is an $\epsilon$-approximate compression scheme with respect to the $\ell_{p}$ loss of size $O\mathopen{}\left(\frac{f(d_{\mathrm{P}})\log(\frac{1}{\epsilon})}{p}\right)$.
@@
Suppose all binary concept classes with VC dimension $\mathrm{VC}$ have a sample compression scheme of size $f(\mathrm{VC})$. Does every $[0,1]$-valued concept class with a finite fat-shattering dimension (at any scale) admit an $\epsilon$-approximate $\ell_{\infty}$ compression scheme of size $O(f(\mathrm{fat}_{c\epsilon})\mathrm{polylog}(\frac{1}{\epsilon},\mathrm{fat}_{c\epsilon}))$ for some $c>0$, where $\mathrm{fat}_{\gamma}$ is the fat-shattering dimension of the concept class at scale $\gamma$?
@@ 4.1 Additional Assumptions: Majority Votes, Proper, and Stable Compression Schemes
Suppose any binary concept class $\mathcal{C}$ with VC dimension $\mathrm{VC}<\infty$ has a compression scheme of size $f(\mathrm{VC})$, which is a proper or a majority vote compression scheme (see Definition 3.4). Then for any $p\in[1,\infty]$, any class with pseudo-dimension $d_{\mathrm{P}}<\infty$ admits an $\epsilon$-approximate compression scheme, with respect to the $\ell_{p}$ loss, of size $O(f(d_{\mathrm{P}}))$.
@@
Suppose any binary concept class $\mathcal{C}$ with VC dimension $\mathrm{VC}<\infty$ has a stable compression scheme of size $f(\mathrm{VC})$. Then, for any $p\in[1,\infty]$, any class with pseudo-dimension $d_{\mathrm{P}}<\infty$ admits an $\epsilon$-approximate compression scheme, with respect to $\ell_{p}$ loss, of size $O\mathopen{}\left(f(d_{\mathrm{P}})\right)$.
@@
Suppose any binary concept class $\mathcal{C}$ with VC dimension $\mathrm{VC}<\infty$ has a compression scheme of size $f(\mathrm{VC})$.
@@
The following theorem slightly deviates from the theme of earlier proofs about reductions to binary compression. Instead, this is a reduction from realizable regression to multiclass classification with infinite labels, by showing that for any concept class $\mathcal{C}\subseteq[0,1]^{\mathcal{X}}$, it holds that $\mathrm{d}_{\mathrm{G}}(\mathcal{C})\leq 4d_{\mathrm{P}}(\mathcal{C})$. Attias et al. [2024] very briefly stated that there is a bounded realizable exact compression scheme for classes with bounded pseudo-dimension, but we provide the full proof in this paper. The proof is in Appendix C.
@@
Suppose any multiclass concept class $\mathcal{C}$ with graph dimension $\mathrm{d}_{\mathrm{G}}<\infty$ has a compression scheme of size $f(\mathrm{d}_{\mathrm{G}})$. Then, any real-valued class with pseudo-dimension $d_{\mathrm{P}}<\infty$ has a compression scheme of size $O(f(4d_{\mathrm{P}}))$. This implies that there is an exact compression scheme of size $O(d_{\mathrm{P}}2^{4d_{\mathrm{P}}})$.
@@ 4.2 Agnostic Approximate Compression for Regression
Suppose that for binary classes with VC dimension $\mathrm{VC}<\infty$, there exists a sample compression scheme of size $f(\mathrm{VC})$. Then, for a real-valued class with pseudo-dimension $d_{\mathrm{P}}<\infty$, there exists an agnostic sample compression scheme with respect to the $\ell_{\infty}$ loss of size $O(f(d_{\mathrm{P}})\log(\frac{1}{\epsilon}))$. Furthermore, for $p\in[1,\infty)$, there exists an agnostic sample compression scheme with respect to the $\ell_{p}$ loss of size $O\mathopen{}\left(\frac{f(d_{\mathrm{P}})\log(\frac{1}{\epsilon})}{p}\right)$.
@@ 5 Compression for Adversarially Robust Classification Against Test-Time Attacks
Let $\mathcal{U}:\mathcal{X}\rightarrow 2^{\mathcal{X}}$ be a perturbation function and let $M=\sup_{x\in\mathcal{X}}|\mathcal{U}(x)|$ be finite. Suppose any binary concept class $\mathcal{C}$ with VC dimension $\mathrm{VC}<\infty$ has a sample compression scheme of size $f(\mathrm{VC})$. Then, $\mathcal{C}$ has an adversarially robust sample compression scheme of size $O(f(\mathrm{VC})\log M)$.
@@
Suppose we have a set $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ that is robustly realizable by a class $\mathcal{C}$. Suppose we have a binary sample compression scheme for this class (with compression and reconstruction functions $\rho_{b},\kappa_{b}$ respectively). We can inflate to $S_{\mathcal{U}}=\{(z,y_{i}):i\in[n],z\in\mathcal{U}(x_{i})\}$ and then apply $\kappa_{b}$ to get a compression set of size $f(\mathrm{VC})$. Suppose $\sup_{x}|\mathcal{U}(x)|=M$ is finite. Then, we can encode each $(z,y_{i})$ using $O(\log M)$ bits by considering the index of $z$ in $\mathcal{U}(x)$, which gives an $O(f(\mathrm{VC})\log M)$ size compression scheme.
@@
Let $\mathcal{U}:\mathcal{X}\rightarrow 2^{\mathcal{X}}$ be a perturbation function and let $M=\sup_{x\in\mathcal{X}}|\mathcal{U}(x)|$ be finite. Suppose any binary concept class $\mathcal{C}$ with VC dimension $\mathrm{VC}<\infty$ has a stable sample compression scheme of size $f(\mathrm{VC})$. Then, $\mathcal{C}$ has an adversarially robust sample compression scheme of size $O(f(\mathrm{VC}))$.
@@
Suppose we have a set $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ that is robustly realizable by a class $\mathcal{C}$. Suppose we have a binary sample compression scheme for this class (with compression and reconstruction functions $\rho_{b},\kappa_{b}$ respectively). We can inflate to $S_{\mathcal{U}}=\{(z,y_{i}):i\in[n],z\in\mathcal{U}(x_{i})\}$ and then apply $\kappa_{b}$ to get a compression set of size $f(\mathrm{VC})$. Suppose $\sup_{x}|\mathcal{U}(x)|=M$ is finite. There is a set $T\subseteq S$ of size $f(\mathrm{VC})$ which contributes to $\kappa_{b}(S_{\mathcal{U}})$. Since the binary compression scheme is stable, we can inflate $T$ to $T_{\mathcal{U}}$ and encode it as $T$, which can be the output of our compression scheme which has size $f(\mathrm{VC})$. For the reconstruction, inflate $T$ to $T_{\mathcal{U}}$ and use $\rho_{b}(\kappa_{b}(T))$ to reconstruct the concept.
@@
Let $\mathcal{C}\subseteq\{0,1\}^{\mathcal{X}}$ be a binary concept class with VC dimension $\mathrm{VC}<\infty$, and let $\mathcal{U}:\mathcal{X}\rightarrow 2^{\mathcal{X}}$ be a perturbation function. Suppose $\mathcal{X}$ has a sample compression scheme of size $f(\mathrm{VC})$. Does there exist an adversarially robust compression scheme of size $O(f(\mathrm{VC}))$ for $\mathcal{C}$?

### v2

@@ 3.1 Additional Assumption: Existence of Proper or Majority Vote Binary Compression
Suppose any binary concept class $\mathcal{C}$ with VC dimension $d_{\mathrm{VC}}<\infty$ has a compression scheme of size $f(d_{\mathrm{VC}})$ which is a proper compression scheme or a majority vote compression scheme (see Definition 3.4). Then any multiclass concept class (label set is allowed to be of infinite size) with a finite graph dimension $d_{\mathrm{G}}<\infty$ admits a compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$.
In particular, we recover the best known bounded-size multiclass compression scheme of size $2^{O\mathopen{}\left(d_{\mathrm{G}}\right)}$ [David et al., 2016], via a reduction to binary compression scheme.
@@
We now construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given realizable $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, let $T=\{((x_{i},y_{i}),1):i\in[n]\}$. We let our compression $\kappa(S)$ return the $(x,y)$ pairs from $\kappa_{b}(T)$, which have size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$. Reconstruction: We have that $\kappa_{b}(T)=\{((x,y),1):(x,y)\in\kappa(S)\}$, so we can immediately recover $\kappa_{b}(T)$. We know that $\rho_{b}(\kappa_{b}(T))$ is correct on all possible $T$. Additionally, since we predict $1$ for at most one $y$ for each $x$, we can conclude that $\rho_{b}(\kappa_{b}(T))$ is correct on all of $S_{\mathcal{Y}}$, as desired. Note that if the reconstruction $\rho_{b}$ is proper, then it will output exactly one $1$ for each $x$. Additionally, it the reconstruction is a majority of learners from the class, where we break ties to favor predicting 00, then it will output at most one $1$ for each $x$.
@@ 3.2 Additional Assumption: Existence of Stable Binary Compression
Suppose that for binary concept classes with finite VC dimension $d_{\mathrm{VC}}<\infty$, there exists a stable sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for multiclass concept classes with a finite graph dimension $d_{\mathrm{G}}<\infty$ and finite label space, there exists a stable sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$.
@@
i.e., it is an inflated $\kappa_{b}(S)$ to include all the labels for each $x$. Since the compression scheme is stable, and since $\kappa_{b}(S_{\mathcal{Y}})\subseteq\kappa_{b}^{\prime}(S_{\mathcal{Y}})\subseteq S_{\mathcal{Y}}$, we must have that $\kappa_{b}(\kappa_{b}^{\prime}(S_{\mathcal{Y}}))=\kappa_{b}(S_{\mathcal{Y}})$. We can let $\kappa(S)$ return the $(x_{i},y_{i})$ pairs such that $x_{i}$ is a member of $\kappa_{b}(S_{\mathcal{Y}})$. This will have size $O(f(d_{\mathrm{G}}))$. Reconstruction: Given $\kappa(S)$, we can reconstruct $\kappa^{\prime}_{b}(S_{\mathcal{Y}})$ by inflating the dataset to include all labels for each $x$. Then, we can apply $\kappa_{b}$ to get $\kappa_{b}(S_{\mathcal{Y}})$ (using the fact that it is a stable compression scheme), and then apply $\rho_{b}$ to get the desired compression scheme.
@@
Consider the class $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$, where $\mathcal{X}=\mathcal{Y}=\mathbb{R}$, of two piecewise thresholds, i.e. $\{g_{t,y_{1},y_{2}}:t,y_{1},y_{2}\in\mathbb{R}\}$ where
@@ 3.3 Agnostic Multiclass Sample Compression Scheme
that for binary classes with VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for a multiclass class with a finite graph dimension $d_{\mathrm{G}}<\infty$ and a finite label set $|\mathcal{Y}|$, there exists an agnostic sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\log|\mathcal{Y}|\right)$.
@@ 4 Compression for Regression
Suppose that for binary classes with VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for a $[0,1]$-real-valued class with pseudo-dimension $d_{\mathrm{P}}<\infty$, there is an $\epsilon$-approximate compression scheme with respect to the $\ell_{\infty}$ loss with size $O(f(d_{\mathrm{P}})\log(\frac{1}{\epsilon}))$. Furthermore, for $p\in[1,\infty)$, there is an $\epsilon$-approximate compression scheme with respect to the $\ell_{p}$ loss of size $O\mathopen{}\left(\frac{f(d_{\mathrm{P}})\log(\frac{1}{\epsilon})}{p}\right)$.
@@
Suppose all binary concept classes with VC dimension $d_{\mathrm{VC}}$ have a sample compression scheme of size $f(d_{\mathrm{VC}})$. Does every $[0,1]$-valued concept class with a finite fat-shattering dimension (at any scale) admit an $\epsilon$-approximate $\ell_{\infty}$ compression scheme of size $O(f(\mathrm{fat}_{c\epsilon})\mathrm{polylog}(\frac{1}{\epsilon},\mathrm{fat}_{c\epsilon}))$ for some $c>0$, where $\mathrm{fat}_{\gamma}$ is the fat-shattering dimension of the concept class at scale $\gamma$?
@@ 4.1 Additional Assumptions: Majority Votes, Proper, and Stable Compression Schemes
Suppose any binary concept class $\mathcal{C}$ with VC dimension $d_{\mathrm{VC}}<\infty$ has a compression scheme of size $f(d_{\mathrm{VC}})$, which is a proper or a majority vote compression scheme (see Definition 3.4). Then for any $p\in[1,\infty]$, any class with pseudo-dimension $d_{\mathrm{P}}<\infty$ admits an $\epsilon$-approximate compression scheme, with respect to the $\ell_{p}$ loss, of size $O(f(d_{\mathrm{P}}))$.
@@
Suppose any binary concept class $\mathcal{C}$ with VC dimension $d_{\mathrm{VC}}<\infty$ has a stable compression scheme of size $f(d_{\mathrm{VC}})$. Then, for any $p\in[1,\infty]$, any class with pseudo-dimension $d_{\mathrm{P}}<\infty$ admits an $\epsilon$-approximate compression scheme, with respect to $\ell_{p}$ loss, of size $O\mathopen{}\left(f(d_{\mathrm{P}})\right)$.
@@
Suppose any binary concept class $\mathcal{C}$ with VC dimension $d_{\mathrm{VC}}<\infty$ has a compression scheme of size $f(d_{\mathrm{VC}})$.
@@
The following theorem slightly deviates from the theme of earlier proofs about reductions to binary compression. Instead, this is a reduction from realizable regression to multiclass classification with infinite labels, by showing that for any concept class $\mathcal{C}\subseteq[0,1]^{\mathcal{X}}$, it holds that $d_{\mathrm{G}}(\mathcal{C})\leq 4d_{\mathrm{P}}(\mathcal{C})$. Attias et al. [2024] very briefly stated that there is a bounded realizable exact compression scheme for classes with bounded pseudo-dimension, but we provide the full proof in this paper. The proof is in Appendix C.
@@
Suppose any multiclass concept class $\mathcal{C}$ with graph dimension $d_{\mathrm{G}}<\infty$ has a compression scheme of size $f(d_{\mathrm{G}})$. Then, any real-valued class with pseudo-dimension $d_{\mathrm{P}}<\infty$ has a compression scheme of size $O(f(4d_{\mathrm{P}}))$. This implies that there is an exact compression scheme of size $O(d_{\mathrm{P}}2^{4d_{\mathrm{P}}})$.
@@ 4.2 Agnostic Approximate Compression for Regression
Suppose that for binary classes with VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for a real-valued class with pseudo-dimension $d_{\mathrm{P}}<\infty$, there exists an agnostic sample compression scheme with respect to the $\ell_{\infty}$ loss of size $O(f(d_{\mathrm{P}})\log(\frac{1}{\epsilon}))$. Furthermore, for $p\in[1,\infty)$, there exists an agnostic sample compression scheme with respect to the $\ell_{p}$ loss of size $O\mathopen{}\left(\frac{f(d_{\mathrm{P}})\log(\frac{1}{\epsilon})}{p}\right)$.
@@ 5 Compression for Adversarially Robust Classification Against Test-Time Attacks
Let $\mathcal{U}:\mathcal{X}\rightarrow 2^{\mathcal{X}}$ be a perturbation function and let $M=\sup_{x\in\mathcal{X}}|\mathcal{U}(x)|$ be finite. Suppose any binary concept class $\mathcal{C}$ with VC dimension $d_{\mathrm{VC}}<\infty$ has a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, $\mathcal{C}$ has an adversarially robust sample compression scheme of size $O(f(d_{\mathrm{VC}})\log M)$.
@@
Suppose we have a set $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ that is robustly realizable by a class $\mathcal{C}$. Suppose we have a binary sample compression scheme for this class (with compression and reconstruction functions $\rho_{b},\kappa_{b}$ respectively). We can inflate to $S_{\mathcal{U}}=\{(z,y_{i}):i\in[n],z\in\mathcal{U}(x_{i})\}$ and then apply $\kappa_{b}$ to get a compression set of size $f(d_{\mathrm{VC}})$. Suppose $\sup_{x}|\mathcal{U}(x)|=M$ is finite. Then, we can encode each $(z,y_{i})$ using $O(\log M)$ bits by considering the index of $z$ in $\mathcal{U}(x)$, which gives an $O(f(d_{\mathrm{VC}})\log M)$ size compression scheme.
@@
Let $\mathcal{U}:\mathcal{X}\rightarrow 2^{\mathcal{X}}$ be a perturbation function and let $M=\sup_{x\in\mathcal{X}}|\mathcal{U}(x)|$ be finite. Suppose any binary concept class $\mathcal{C}$ with VC dimension $d_{\mathrm{VC}}<\infty$ has a stable sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, $\mathcal{C}$ has an adversarially robust sample compression scheme of size $O(f(d_{\mathrm{VC}}))$.
@@
Suppose we have a set $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ that is robustly realizable by a class $\mathcal{C}$. Suppose we have a binary sample compression scheme for this class (with compression and reconstruction functions $\rho_{b},\kappa_{b}$ respectively). We can inflate to $S_{\mathcal{U}}=\{(z,y_{i}):i\in[n],z\in\mathcal{U}(x_{i})\}$ and then apply $\kappa_{b}$ to get a compression set of size $f(d_{\mathrm{VC}})$. Suppose $\sup_{x}|\mathcal{U}(x)|=M$ is finite. There is a set $T\subseteq S$ of size $f(d_{\mathrm{VC}})$ which contributes to $\kappa_{b}(S_{\mathcal{U}})$. Since the binary compression scheme is stable, we can inflate $T$ to $T_{\mathcal{U}}$ and encode it as $T$, which can be the output of our compression scheme which has size $f(d_{\mathrm{VC}})$. For the reconstruction, inflate $T$ to $T_{\mathcal{U}}$ and use $\rho_{b}(\kappa_{b}(T))$ to reconstruct the concept.
@@
Let $\mathcal{C}\subseteq\{0,1\}^{\mathcal{X}}$ be a binary concept class with VC dimension $d_{\mathrm{VC}}<\infty$, and let $\mathcal{U}:\mathcal{X}\rightarrow 2^{\mathcal{X}}$ be a perturbation function. Suppose $\mathcal{X}$ has a sample compression scheme of size $f(d_{\mathrm{VC}})$. Does there exist an adversarially robust compression scheme of size $O(f(d_{\mathrm{VC}}))$ for $\mathcal{C}$?